# Efficient Measure of Information Backflow with Quasi-stochastic Process

### v1

@@ Abstract
Characterization and quantification of non-Markovian dynamics in open quantum systems is a topical issue in the rapidly developing field of quantum computation and quantum communication. A standard approach based on the notion of information backflow detects the flow of information from the environment back to the system. Numerous measures of information backflow have been proposed using different definitions of distinguishability between pairs of quantum states. These measures, however, necessitate optimization over the state space which can be analytically challenging or numerically demanding. Here, we propose an alternative witness and measure of information backflow that is explicitly state-independent by utilizing the concept of quasiprobability representation and recent advances in the theory of majorization for quasiprobabilities. We illustrate its use over several paradigmatic examples, demonstrating consistent Markovian conditions with known results and also reported necessary and sufficient condition for qutrit system in random unitary channel. The paper concludes with discussions on the foundational implications of quantum dynamical evolution.
Introduction—In recent years, the study of open quantum systems beyond the standard Markovian approximation has gained considerable interest. In Markovian dynamics, information in a system dissipates continuously to the environment. Interestingly, it has been observed that non-Markovian dynamics exhibit a memory effect such that information can flow back to the system in a revival of the quantum features [1, 2, 3]. This phenomenon has become a subject of interest in the study of decoherence [4, 5], quantum thermodynamics [6], and quantum technologies in general [7]. Thus, an efficient tool to identify and quantify non-Markovianity features is highly desirable.
There have been many proposals for a witness of non-Markovianity with the central object being the quantum dynamical map, $\Lambda_{t}$, $t\geq 0$, which encapsulates the evolution of the open system. In the standard Hilbert space formalism, they are described by completely positive and trace preserving (CPTP) maps that can inform the evolution of a quantum state at any time $t$ via $\rho(t)=\Lambda_{t}[\rho(0)]$. To track information backflow, one is only required to use some notion of distinguishability between two initial states and track its evolution over time. The most natural choice is the trace distance between two states defined as
@@
where the LHS is a monotonically decreasing function for all time $t\geq 0$ if $\Lambda_{t}$ is Markovian. Therefore, following [8], the BLP (Breuer-Laine-Piilo) measure defined as
[…]
@@
with $\sigma(\rho_{1},\rho_{2};t)\equiv\frac{d}{dt}D(\Lambda_{t}[\rho_{1}],\Lambda_{t}[\rho_{2}])$, quantifies the degree of memory effects. The integral is taken only whenever the information flow $\sigma(\rho_{1},\rho_{2};t)$ is positive — an indication of violating Markovianity. As non-Markovianity is a feature solely due to the dynamics, maximization over all pair states is taken to remove the state dependence.
@@
The above is not the only existing measure of information backflow. Many other proposals utilizing alternative or more general definitions of distances, such as Helstrom matrix [9, 10], and also entropic formulation [11, 12] have been carried out. They all, however, suffer from exact computation as optimizing over state space can be challenging, especially for higher dimensional systems. Even though progress has been made to simplify the optimization task [13], it would be ideal to completely characterize and measure non-Markovianity without this complication. It is also noteworthy to point out that there exists another notion of non-Markovianity based on divisibility [14, 15]. P- or CP-divisibility of a dynamical map concerns whether the intermediate map is still positive or completely positive, respectively, which is then associated with being Markovian. The hierarchical relation between divisibility and information backflow has been studied extensively with the former constituting a stricter criterion for non-Markovianity. Here, we shall only focus on the latter.
@@
In this Letter, we propose a novel measure of non-Markovianity, in particular, information backflow, that requires no optimization component using the framework of quasiprobability representation (QPR) [16, 17]. QPR has been gaining popularity as it provides an alternative mathematical formalism of quantum theories that are reminiscent of probability theory but with the appearance of negative values — widely attributed as the signature of nonclassicality [18, 19, 20, 21, 22, 23]. As such, it has been particularly useful in the study of classical versus quantum theory, identifying resources for quantum advantage, and in general provides new insights about quantum theories. Following the latter point, we will further comment on the implication of our non-Markovianity witness in a quantum foundational perspective. We note that previous work [24] had considered the use of frame representation to develop a new measure. However, similar to the previously mentioned measures, it also involves maximizing over all possible pairs of initial states. Crucially, their work focuses on studying non-Markovianity in continuous variable system, but here we focus on the discrete case.
@@
Quasiprobability representation—In QPR, elements of quantum theories, such as states and unitaries, are described by vectors or matrices consisting of real values that typically obey a normalization constraint. To translate quantum objects defined in the standard Hilbert space formalism into its QPR, the mapping is achieved by the so-called frame and its dual frame [16, 17]. Hence, it is also colloquially known as frame representation. In the following, we will introduce the mathematical formalism of QPR.
@@
For a system in Hilbert space $\mathcal{H}^{d}$, a frame $\{F_{j}\}$ is defined as a set of operators that span the Hermitian space. That is, it forms, in general, an overcomplete basis for the space of all Hermitian operators acting on $\mathcal{H}^{d}$. For a given frame, the dual frame $\{G_{j}\}$ is also a complete basis of a Hermitian space that satisfy $A=\sum_{j}\text{Tr}\left[AF_{j}\right]G_{j}$ for all Hermitian operator $A$. Note that the dual frame is generally not unique except for the case when the frame is minimal, i.e., |$|\{F_{j}\}|=d^{2}$. In this case, they obey the relation $\text{Tr}\left[F_{j}G_{k}\right]=\delta_{jk}$. The QPR of a state $\rho$, channel $\mathcal{E}$, and measurement effect $M$ are then obtained via
@@
Note that the above always return real values and they correspond to the entries of the column vector $\mathbf{q}^{\rho}=[q^{\rho}_{j}]$, square matrix $S^{\mathcal{E}}=[S^{\mathcal{E}}_{jk}]$, and row vector $\mathbf{v}^{M}=[v^{M}_{j}]$. From here, the Born rule probability of measuring a state undergoing an evolution beforehand is simply calculated following vector/matrix operation:
@@
Here, we consider a normalization condition on the frame operators, $\sum_{j}F_{j}=\mathbb{1}$. This implies that $\sum_{j}q^{\rho}_{j}=1$ so it can interpreted as a quasiprobability distribution. Moreover, we also imposed that $\text{Tr}\left[G_{j}\right]=1\,\forall\,j$ to ensure that summing up the effects in a POVM (positive operator-value measure) returns a unit measurement: $\sum_{k}v^{M_{k}}_{j}=1\,\forall j$, where $\{M_{k}|M_{k}\geq 0,\sum_{k}M_{k}=\mathbb{1}\}$ is a POVM. As a result, $v^{M_{k}}_{j}$ can be thought of as a conditional quasiprobability distribution. Lastly, based on the above condition, one also finds that $\sum_{j}S^{\mathcal{E}}_{jk}=1\,\forall k$, which means that it is now a quasi-stochastic matrix. If the channel is unital, i.e., $\mathcal{E}[\mathbb{1}]=\mathbb{1}$, then $S^{\mathcal{E}}$ is quasi-bistochastic, [^1] In some literature, ‘bistochastic’ is also referred as ‘doubly stochastic’. $\sum_{j}S^{\mathcal{E}}_{jk}=\sum_{k}S^{\mathcal{E}}_{jk}=1$. As can be seen, the elements of quantum theories have been described in a language that is similar to classical theories and hence is natural to use for the study of classical-quantum distinction.
Majorization theory for quasiprobability distributions—To derive the main results of this paper, we shall make use of the observations made in [25]. There, the authors extended the theory of majorization for quasiprobabilities and shown that certain family of entropries are well-defined and meaningful even for quasiprobabilities. Of particular interest, Theorem 9 of [25] states that:
If $\alpha=\frac{2a}{2b-1}$ for positive integers $a,b$ with $a\geq b$, then $H_{\alpha}(\mathbf{q})$ is well-defined on the set of quasi-distributions, and if $\mathbf{q}\succ\mathbf{q}^{\prime}$ for two quasi-distributions $\mathbf{q},\mathbf{q}^{\prime}$ then $H_{\alpha}(\mathbf{q})\leq H_{\alpha}(\mathbf{q}^{\prime})$.
Here, $H_{\alpha}(\mathbf{q})$ is the Renyi-$\alpha$ entropy [26] defined as
$H_{\alpha}(\mathbf{q})=\frac{1}{1-\alpha}\log\left[\sum_{k}q_{k}^{\alpha}\right]\,,$ (8)
@@
The result above establishes the Schur-concavity property and monotonicity of Renyi-$\alpha$ entropy on the set of quasiprobability distributions under a subset of $\alpha$ values. Moreover, the choice of $\alpha$ also ensures that the function is real-valued and satisfies many of the properties that Renyi originally invoked [27].
Main results—In the following, we shall take advantage of the result above to derive a new witness and measure of non-Markovianity. Let us make the choice of $\alpha=2$ for the Renyi-$\alpha$ entropy. This form is also known as collision entropy, and its usage has been explored previously as entropy of Wigner function [28], in relation to uncertainty principle [29], and in axiomatization of quantum theories [30, 31, 32].
Considering state transforming over time given by $\mathbf{q}(t)=S^{\Lambda_{t}}\mathbf{q}(0)$, where we denote $\mathbf{q}\equiv\mathbf{q}^{\rho}$ for brevity, then it follows from the above that
@@
if $S^{\Lambda_{t}}$ is bistochastic for all time $t\geq 0$. One can then immediately show that under Markovian condition the monotonicity holds, and violation occurs for non-Markovian case. This formulation, however, is akin to the entropic-based formulation in [11] and would still entail a maximization over the state space that is inefficient.
Surprisingly, we show that it is possible to further simplify the above relation and remove the dependence on state. Notice that the Renyi-2 entropy can be expressed more compactly as
@@
where $\mathbf{q}^{T}$ is the transpose of $\mathbf{q}$ into a row vector. Therefore, Eq. (9) can be written as $-\log[(S^{\Lambda_{t}}\mathbf{q})^{T}(S^{\Lambda_{t}}\mathbf{q})]=-\log[\mathbf{q}^{T}(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}\mathbf{q}]\geq-\log\mathbf{q}^{T}\mathbf{q}$ and can be simplified neatly as
@@
Under Markovian condition, the quasi-stochastic map of a channel $\Lambda_{t}$ must then obey this monotonic relation. More explicitly, the eigenvalues of $(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}$ must be a monotonically decreasing function over time, and the violation of the monotonicity signifies non-Markovianity. As such, in the same spirit as previously proposed measures, we define our measure of non-Markovianity as
@@
We remark that this formulation is reminiscent of the measure for P- or CP-divisibility [14] such that it is a sole function on the dynamical map so no optimization over state is required and the eigenvalues of the map obey a monotonic relation.
Let us now discuss some of the properties of the above measure. Firstly, we have that $(S^{\Lambda_{t}})^{T}=S^{\Lambda_{t}^{\dagger}}$. From here, it is easy to see that the equality in Eq. (11) is saturated if the channel is unitary, which reduces into orthogonality relation. This is obvious from the basic property of unitary matrices: $U^{\dagger}=U^{-1}$. Interestingly, it has been shown before that an axiom based on the orthogonality of the quasi-bistochastic map enable for the reconstruction of qubit and its dynamics [32].
Secondly, an equivalent witness can be obtained by swapping the matrices in Eq. (11), i.e, $S^{\Lambda_{t}}(S^{\Lambda_{t}})^{T}\leq\mathbb{1}$. This is clear as the eigenvalues of $AB$ is equal to $BA$. Moreover, both instances of the product matrix yield a symmetric matrix. The observation on the symmetry of the witness has a foundational implications that will be discussed later.
Next, we shall see our measure of non-Markovianity in action over several paradigmatic examples with qubit and qutrit systems. Notice that the measure above is derived without any reference for a specific frame representation. Thus, the result works in general for any choice of quasiprobability representation. In order to make specific calculation, here we pick the discrete Wigner representation with Wootters’ representation for the qubit case [33] and Gross’ for the qudit case (odd prime dimension) [34]. In this representation, they belong to the family of normal quasiprobability representation as their frame and dual frame operators are proportional to each other [35]. In fact, they are related via $G_{j}=dF_{j}$ since the frame is minimal. The Wootters’ discrete Wigner representation has frame operators given by

### v2

# Efficient measure of information backflow with a quasistochastic process

@@ Abstract
Characterization and quantification of non-Markovian dynamics in open quantum systems are topical issues in the rapidly developing field of quantum computation and quantum communication. A standard approach based on the notion of information backflow detects the flow of information from the environment back to the system. Numerous measures of information backflow have been proposed using different definitions of distinguishability between pairs of quantum states. These measures, however, necessitate optimization over the state space, which can be analytically challenging or numerically demanding. Here we propose an alternative witness and measure of information backflow that is explicitly state independent by utilizing the concept of quasiprobability representation and recent advances in the theory of majorization for quasiprobabilities. We illustrate its use over several paradigmatic examples, demonstrating consistent Markovian conditions with known results and also reported necessary and sufficient conditions for the qutrit system in a random unitary channel. The paper concludes with a discussion of the foundational implications of quantum dynamical evolution.
I Introduction
In recent years, the study of open quantum systems beyond the standard Markovian approximation has gained considerable interest. In Markovian dynamics, information in a system dissipates continuously to the environment. Interestingly, it has been observed that non-Markovian dynamics exhibit a memory effect such that information can flow back to the system in a revival of the quantum features [1, 2, 3]. This phenomenon has become a subject of interest in the study of decoherence [4, 5], quantum thermodynamics [6], and quantum technologies in general [7]. Thus, an efficient tool to identify and quantify non-Markovianity features is highly desirable.
There have been many proposals for a witness of non-Markovianity with the central object being the quantum dynamical map $\Lambda_{t}$, $t\geq 0$, which encapsulates the evolution of the open system. In the standard Hilbert space formalism, the maps are described as completely positive and trace-preserving (CPTP) maps that can inform the evolution of a quantum state at any time $t$ via $\rho(t)=\Lambda_{t}[\rho(0)]$. To track information backflow, one is only required to use some notion of distinguishability between two initial states and track its evolution over time. The most natural choice is the trace distance between two states defined as
@@ I Introduction
where the left-hand side is a monotonically decreasing function for all time $t\geq 0$. When monotonicity holds at all times, we say that $\Lambda_{t}$ is Markovian. In contrast, there exists quantum dynamics that breaks the monotonicity for some time interval, which results in a temporary revival of quantum information into the system. In such cases, we say that $\Lambda_{t}$ is non-Markovian. Therefore, following [8], the Breuer-Laine-Piilo (BLP) measure defined as
$\mathcal{N}_{\text{BLP}}=\max_{\rho_{1},\rho_{2}}\int_{\sigma>0}dt\,\sigma(\rho_{1},\rho_{2};t)\,$, (3)
@@
with $\sigma(\rho_{1},\rho_{2};t)\equiv\frac{d}{dt}D(\Lambda_{t}[\rho_{1}],\Lambda_{t}[\rho_{2}])$, quantifies the degree of memory effects. The integral is taken only whenever the information flow $\sigma(\rho_{1},\rho_{2};t)$ is positive, an indication of violating Markovianity. As non-Markovianity is a feature solely due to the dynamics, maximization over all pair states is taken to remove the state dependence.
@@
The above is not the only existing measure of information backflow. Many other proposals utilizing alternative or more general definitions of distances, such as the Helstrom matrix [9, 10], and also entropic formulation [11, 12] have been carried out. To our knowledge, these measures suffer from exact computation as optimizing over state space can be challenging, especially for higher-dimensional systems. Even though progress has been made to simplify the optimization task [13], it would be ideal to completely characterize and measure non-Markovianity without this complication. It is also noteworthy that there exists another notion of non-Markovianity based on divisibility [14, 15]. P- or CP-divisibility of a dynamical map concerns whether the intermediate map is still positive or completely positive, respectively, which is then associated with being Markovian. The hierarchical relation between divisibility and information backflow has been studied extensively, with the former constituting a stricter criterion for non-Markovianity. Unlike the distinguishability-based measure, measures of non-Markovianity based on indivisibility are free of optimization. Hence, in this work, we focus on information backflow as our notion of non-Markovianity and find an efficient measure for it.
@@
In this paper, we propose a novel measure of non-Markovianity, in particular, information backflow, that requires no optimization component using the framework of a quasiprobability representation (QPR) [16, 17]. The QPR has been gaining popularity as it provides an alternative mathematical formalism of quantum theories that is reminiscent of probability theory but with the appearance of negative values, widely attributed as the signature of nonclassicality [18, 19, 20, 21, 22, 23]. As such, it has been particularly useful in the study of classical versus quantum theory, identifying resources for quantum advantage, and in general providing new insights about quantum theories. Following the latter point, we will further comment on the implication of our non-Markovianity witness from a quantum foundational perspective. We note that previous work [24] had considered the use of frame representation to develop a new measure. However, similar to the previously mentioned measures, it also involves maximizing over all possible pairs of initial states. Crucially, that work focused on studying non-Markovianity in continuous-variable systems, but here we focus on the discrete case.
II Quasiprobability representation
@@
In the QPR, elements of quantum theories, such as states and unitaries, are described by vectors or matrices consisting of real values that typically obey a normalization constraint. To translate quantum objects defined in the standard Hilbert space formalism into its QPR, the mapping is achieved by the so-called frame and its dual frame [16, 17]. Hence, it is also colloquially known as frame representation. In the following, we will introduce the mathematical formalism of QPR.
@@
For a system in Hilbert space $\mathcal{H}^{d}$, a frame $\{F_{j}\}$ is defined as a set of operators that span the Hermitian space, that is, it forms in general an overcomplete basis for the space of all Hermitian operators acting on $\mathcal{H}^{d}$. For a given frame, the dual frame $\{G_{j}\}$ is also a complete basis of a Hermitian space that satisfies $A=\sum_{j}\text{Tr}\left[AF_{j}\right]G_{j}$ for all Hermitian operator $A$. Note that the dual frame is generally not unique except for the case when the frame is minimal, i.e., $|\{F_{j}\}|=d^{2}$. In this case, the frames obey the relation $\text{Tr}\left[F_{j}G_{k}\right]=\delta_{jk}$. The QPR of a state $\rho$, channel $\mathcal{E}$, and measurement effect $M$ is then obtained via
@@
Note that the above always return real values and they correspond to the entries of the column vector $\mathbf{q}^{\rho}=[q^{\rho}_{j}]$, square matrix $S^{\mathcal{E}}=[S^{\mathcal{E}}_{jk}]$ (with row index $j$ and column index $k$), and row vector $\mathbf{v}^{M}=[v^{M}_{j}]$. From here, the Born rule probability of measuring a state undergoing an evolution beforehand is simply calculated following vector-matrix operation:
@@
Here, we consider a normalization condition for the frame operators, $\sum_{j}F_{j}=\mathbb{1}$. This implies that $\sum_{j}q^{\rho}_{j}=1$, so $\mathbf{q}^{\rho}$ can be interpreted as a quasiprobability distribution. Moreover, we also impose that $\text{Tr}\left[G_{j}\right]=1\,\forall\,j$ to ensure that summing up the effects in a positive-operator-valued measure (POVM) returns a unit measurement $\sum_{k}v^{M_{k}}_{j}=1\,\forall\,j$, where $\{M_{k}\,|\,M_{k}\geq 0,\sum_{k}M_{k}=\mathbb{1}\}$ is a POVM. As a result, $v^{M_{k}}_{j}$ can be thought of as a conditional quasiprobability distribution. Finally, based on the above condition, we also find that $\sum_{j}S^{\mathcal{E}}_{jk}=1\,\forall\,k$, which means that it is now a quasistochastic matrix. If the channel is unital, i.e., $\mathcal{E}[\mathbb{1}]=\mathbb{1}$, then $S^{\mathcal{E}}$ is quasi-bistochastic, [^1] In some literature, “bistochastic” is also referred as “doubly stochastic.” $\sum_{j}S^{\mathcal{E}}_{jk}=\sum_{k}S^{\mathcal{E}}_{jk}=1$. As can be seen, the elements of quantum theories have been described in a language that is similar to classical theories and hence is natural to use for the study of classical-quantum distinction.
III Majorization of quasiprobability distributions
To derive the main results of this paper, we make use of the observations made in [25]. There the authors extended the theory of majorization for quasiprobabilities and showed that a certain family of entropies is well defined and meaningful even for quasiprobabilities. Of particular interest, Theorem 9 of [25] states the following: “If $\alpha=\frac{2a}{2b-1}$ for positive integers $a,b$ with $a\geq b$, then $H_{\alpha}(\mathbf{q})$ is well defined on the set of quasidistributions, and if $\mathbf{q}\succ\mathbf{q}^{\prime}$ for two quasidistributions $\mathbf{q},\mathbf{q}^{\prime}$ then $H_{\alpha}(\mathbf{q})\leq H_{\alpha}(\mathbf{q}^{\prime})$.” Here $H_{\alpha}(\mathbf{q})$ is the Rényi-$\alpha$ entropy [26] defined as
$H_{\alpha}(\mathbf{q})=\frac{1}{1-\alpha}\log_{2}\left[\sum_{k}q_{k}^{\alpha}\right]$ (8)
@@ III Majorization of quasiprobability distributions
The result above establishes the Schur-concavity property and monotonicity of Rényi-$\alpha$ entropy on the set of quasiprobability distributions under a subset of $\alpha$ values. Moreover, the choice of $\alpha$ also ensures that the function is real-valued and satisfies many of the properties that Rényi originally invoked [27].
IV Main results
In the following, we take advantage of the result above to derive a new witness and measure of non-Markovianity. We make the choice of $\alpha=2$ for the Rényi-$\alpha$ entropy. This form is also known as collision entropy, and its usage has been explored previously as the entropy of the Wigner function [28], in relation to the uncertainty principle [29], and in axiomatization of quantum theories [30, 31, 32].
Considering a state transforming over time given by $\mathbf{q}(t)=S^{\Lambda_{t}}\mathbf{q}(0)$, where we define $\mathbf{q}\equiv\mathbf{q}^{\rho}$ for brevity, it follows from the above that
@@ IV Main results
if $S^{\Lambda_{t}}$ is bistochastic for all time $t\geq 0$. One can think of the Rényi-2 entropy of the quantum state as measuring the disorder of the quantum state in the usual manner. States with high disorder (low information content) will have higher entropy than states with low disorder (high information content). Following this interpretation, Markovian dynamics will result in a monotonic loss of information to the environment and violation of the monotonicity condition indicates memory effects (information backflow). Our formulation here is distinct from the usual approach based on monotonicity of the trace distance and is more akin to the entropic-based formulation in [11].
In the current form, one can define a measure of non-Markovianity but it would still entail a maximization over state. Surprisingly, we show that it is possible to further simplify the above relation and remove the dependence on state. Note that the Rényi-2 entropy can be expressed more compactly as
@@
where $\mathbf{q}^{T}$ is the transpose of $\mathbf{q}$ into a row vector. Therefore, Eq. (9) can be written as $\log[(S^{\Lambda_{t}}\mathbf{q})^{T}(S^{\Lambda_{t}}\mathbf{q})]=-\log[\mathbf{q}^{T}(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}\mathbf{q}]\geq-\log\mathbf{q}^{T}\mathbf{q}$ and can be simplified neatly as
@@
Under Markovian condition, the quasi-stochastic map of a channel $\Lambda_{t}$ must then obey this monotonic relation. More explicitly, the eigenvalues of $(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}$ must be a monotonically decreasing function over time, and the violation of the monotonicity signifies non-Markovianity. As such, we define our measure of non-Markovianity as
$\mathcal{N}$ $=$ $\int_{\zeta>0}dt\,\zeta(t)$
The measure defined here is similar in spirit to previously proposed measures such that they sum up over every instance when non-Markovianity is observed based on the respective witness but will yield different values in general. This will be more evident in the examples below. Finally, we remark that this formulation is reminiscent of the measure for P- or CP-divisibility [14] such that it is a sole function on the dynamical map so no optimization over state is required and the eigenvalues of the map obey a monotonic relation.
Let us now discuss some of the properties of the above measure. First, we have that $(S^{\Lambda_{t}})^{T}=S^{\Lambda_{t}^{\dagger}}$. From here it is easy to see that the equality in Eq. (11) is saturated if the channel is unitary, which reduces to a orthogonality relation. This is obvious from the basic property of unitary matrices: $U^{\dagger}=U^{-1}$. Interestingly, it has been shown before that an axiom based on the orthogonality of the quasibistochastic map enables the reconstruction of the qubit and its dynamics [32].
Second, an equivalent witness can be obtained by swapping the matrices in Eq. (11), i.e., $S^{\Lambda_{t}}(S^{\Lambda_{t}})^{T}\leq\mathbb{1}$. This is clear as the eigenvalue of $AB$ is equal to $BA$. Moreover, both instances of the product matrix yield a symmetric matrix. The observation of the symmetry of the witness has a foundational implication that will be discussed later.
@@
V Examples
Next we will see our measure of non-Markovianity in action over several paradigmatic examples with qubit and qutrit systems. Notice that the measure above is derived without any reference for a specific frame representation. Thus, the result works in general for any choice of quasiprobability representation. In order to make a specific calculation, here we pick the discrete Wigner representation with Wootters’ representation for the qubit case [34] and Gross’s for the qudit case (odd prime dimension) [35]. In this representation, they belong to the family of normal quasiprobability representations as their frame and dual frame operators are proportional to each other [36]. In fact, they are related via $G_{j}=dF_{j}$ since the frame is minimal. Wootters’ discrete Wigner representation has frame operators given by